# Benchmarking Multi-Modal Retrieval for Long Documents

### v1

@@ 3.4. Statistics and Analysis
Document Analysis. As shown in Table 3, MMDocIR evaluation set includes 313 long documents averaging 65.1 pages, categorized into ten main domains. Different domains feature distinct distributions of multi-modal information. For instance, research reports, tutorials, workshops, and brochures predominantly contain images, whereas financial and industry documents are table-rich. In contrast, government and legal documents primarily comprise text. Overall, the modality distribution is: Text (60.4%), Image (18.8%), Table (16.7%), and other modalities (4.1%).
@@
Question and Annotation Analysis. MMDocIR encompasses 1,658 questions, 2,107 page labels, and 2,638 layout labels. The modalities required to answer these questions distribute across four categories: Text (44.7%), Image (21.7%), Table (37.4%), and Layout/Meta (11.5%). The “Layout/Meta” category encompasses questions related to layout information and meta-data statistics. Notably, the dataset poses several challenges: 254 questions necessitate cross-modal understanding, 313 questions demand evidence across multiple pages, and 637 questions require reasoning based on multiple layouts. These complexities highlight the need for advanced multi-modal reasoning and contextual understanding.
4. MMDocIR: Training Set
4.1. Document Corpus Collection
The training set corpora are mainly collected from DocVQA benchmarks, adhering to similar selection criteria (see Section 3.1).
  • MP-DocVQA (Tito et al., 2023) contains 47,952 images collected from Industry Documents Library (IDL)  [^5] https://www.industrydocuments.ucsf.edu/. We group the 47,952 document images into separate document files, and obtain 875 long documents (46.8 pages on average) with 15,266 QA pairs.
  • SlideVQA (Tanaka et al., 2023) contains 2,619 slide documents collected from slideshare  [^6] https://www.slideshare.net/ and covering 39 topics. Note that SlideVQA contains only the first 20 pages of each slide decks. In our research, we manually collect the complete slide decks, and obtain 2,011 long documents (averaging 49.3 pages) with 11,066 QA pairs.
  • TAT-DQA (Zhu et al., 2022) consists of 3,067 document pages from financial reports  [^7] https://www.annualreports.com/, dated between 2018 and 2020. Note that neither original documents nor links is provided. We use OCR to extract text in the pages, and use search engine to find relevant documents. After careful tracing and recognition, we identify 163 original documents (averaging 147.3 pages) with 15,814 QA pairs.
  • ArXivQA (Li et al., 2024) comprises 32k figures cropped from academic pages  [^8] https://arxiv.org/. We use the arXiv DOIs provided to collect the academic papers. After careful tracing, recognition, and document length filtering, we identify 1,579 documents averaging 18.4 pages.
  • SciQAG (Wan et al., 2024) consists of 22,728 papers in 24 scientific disciplines, collected from Web of Science (WoS) Core Collection database. We sample 50 documents from each discipline, and manually collect 1,197 papers using the DOIs provided.
  • DUDE (Landeghem et al., 2023) provides 5,019 documents from aggregate websites9991: archive.org, 2: http://commons.wikimedia.org/, 3: http://documentcloud.org/. It covers a broad range of domains, including medical, legal, technical, and financial, among others, to evaluate models’ ability to handle diverse topics and the specific knowledge each requires. We filter out short documents and obtain 779 relatively long documents (averaging 15.6 pages) with 3,173 QA pairs.
  • CUAD (Hendrycks et al., 2021) provides 510 commercial legal contracts, collected from Electronic Data Gathering, Analysis, and Retrieval (EDGAR) [^10] https://www.sec.gov/search-filings. We filter out short documents in CUAD and obtain 274 long documents (29.6 pages on average) with 11,234 QA pairs.
4.2. Label Construction
The page labels can be converted for MP-DocVQA, SlideVQA, and DUDE datasets. Among them, only DUDE provides layout labels.
SciQAG provides only question and answer in texts. We utilize them to reversely obtain the page-level and layout-level labels. Specifically, we first use MinerU to obtain layout-level passage chunks. For each QA pair, we deploy E5 and BGE retrievers to obtain question-passage and answer-passage similarity scores against all extracted passage chunks. If both scores rank within top 3 for a certain passage chunk, we confirm this layout to be the layout-level labels for the given QA pair.
Similarly, ArXivQA provides only cropped images, without document page/layout labels. We first use MinerU to obtain layout-level image. For each cropped image, we calculate the image similarity between all extracted images, and keep only the highest scoring one. Subsequently, we manually examine if the selected image matches the cropped image. In this way, we filter around 20% unmatched images and their QA pairs, resulting 1,579 QA pairs with page-level and layout-level labels.
For TAT-DQA, the layout-level labels are provided for sampled single page. We need to localize the page index of the sampled page. We utilize PDF mapping technique to retrieve the best matched page for every sampled page in this document. Then, we manual examine if the retrieved page is identical to the given page, and correct the labels if there were any errors.
The overall statistics (e.g., document information, modality distribution, domain, etc) of MMDocIR training set are in Table 3.3.
@@ 4.2. Label Construction
lll—cccccccccc—cc

### v2

@@ 3.5 Statistics and Analysis
Document Analysis. As shown in Table 3, MMDocIR evaluation set includes 313 long documents, averaging 65.1 pages, categorized into 10 domains. Different domains feature distinct multimodal distribution. The overall modality distribution is as follows: text (60.4%), image (18.8%), table (16.7%), and others (4.1%), with fine-grained distribution shown in Figure 4(a).
@@
Question and Annotation Analysis. MMDocIR includes 1,658 questions, and 2,107 page and 2,638 layout labels. The evidence spans 4 modalities: text (44.7%), image (21.7%), table (37.4%), and layout/meta (11.5%). Notably, MMDocIR presents several challenges: 254 questions require cross-modal understanding, 313 questions require evidence across multiple pages, and 637 questions require reasoning over multiple layouts.
4 MMDocIR: Training Set
4.1 Document Corpus Collection
After screening related DocVQA datasets, we collect our training set corpora from 7 datasets, namely MP-DocVQA Tito et al. (2023), SlideVQA Tanaka et al. (2023), TAT-DQA Zhu et al. (2022), SciQAG Wan et al. (2024), DUDE Landeghem et al. (2023), and CUAD Hendrycks et al. (2021). Since most of these datasets do not provide original document, we invest significant efforts in tracing and recovering the original document, as detailed in Appendix B.4.
4.2 Label Construction and statistics
We use semi-automated construction pipeline to generate page-level and layout-level labels for datasets that lack them, referring to Appendix B.5 and B.6 for more details of construction process. Notably, layout annotations are missing from most existing datasets, as we manage to obtain or construct layout-level labels for only 4 datasets. The overall statistics (e.g., document information, modality distribution, domain, etc) of MMDocIR training set are summarized in Table 4.
@@ 4.2 Label Construction and statistics
lll|cccccccccc|cc
Report &Indu. Worksh. Paper ure Report book ment Macro Micro
Recall@$k=1$
VLM-text
DPR 32.3 25.5 27.0 31.0 28.4 18.8 23.5 31.2 38.3 16.1 27.2 26.9
ColBERT 48.6 42.8 51.1 46.2 36.0 36.8 49.6 60.9 59.5 26.3 45.8 44.9
BGE 48.8 30.9 47.1 40.8 37.6 28.4 43.4 51.9 48.9 28.5 40.6 39.6
E5 48.1 30.0 50.4 39.4 41.1 29.7 40.9 52.8 51.1 24.1 40.8 39.5
Contriever 45.5 31.2 49.8 41.5 39.4 29.4 45.2 55.3 51.1 20.4 40.9 39.7
GTE 46.5 26.3 48.7 38.9 35.9 27.0 46.2 50.1 45.8 24.1 38.9 37.9
[…]
Image
DSEwiki-ss 53.0 50.0 54.0 48.7 45.1 43.0 51.5 46.9 54.2 33.6 48.0 47.5
DSEdocmatix 52.3 40.4 56.1 51.7 45.8 43.5 53.8 53.7 58.3 46.7 50.2 50.1
ColPali 56.0 51.8 58.6 55.9 52.0 47.2 57.9 53.9 64.0 32.8 53.0 52.7
Recall@$k=3$
VLM-text
DPR 52.2 44.2 43.5 54.6 52.0 35.1 44.4 53.9 57.2 25.5 46.3 46.2
ColBERT 70.1 64.4 70.3 72.3 59.1 55.3 71.1 81.3 70.8 34.3 64.9 64.8
BGE 71.5 48.2 68.8 65.7 56.2 46.5 66.1 69.9 72.0 32.1 59.7 59.6
E5 68.4 45.7 68.1 63.7 60.1 44.0 69.3 72.3 78.8 32.8 60.3 59.3
Contriever 69.4 55.3 68.3 64.9 56.9 46.2 69.9 71.1 72.0 32.1 60.6 59.7
GTE 71.1 44.5 67.2 64.4 54.3 43.0 70.6 71.9 68.2 31.4 58.7 58.3
[…]
Image
DSEwiki-ss 75.4 65.0 73.9 79.8 69.5 63.5 75.4 71.5 81.4 50.4 70.6 71.4
DSEdocmatix 75.4 67.5 73.3 80.0 66.3 61.6 72.8 76.4 82.6 57.7 71.4 71.8
ColPali 77.6 71.8 79.4 83.4 72.6 66.1 80.0 80.4 86.4 49.6 74.7 75.0
Recall@$k=5$
VLM-text
DPR 66.5 60.1 56.0 68.9 58.8 43.8 57.1 68.6 64.8 33.6 57.8 57.8
ColBERT 78.8 74.0 78.7 82.3 66.1 60.8 77.0 88.5 78.0 38.7 72.3 72.3
BGE 79.5 65.8 71.3 76.8 62.4 56.0 77.2 77.4 79.5 38.0 68.4 68.5
E5 76.9 64.2 75.3 74.4 67.4 52.0 78.5 78.6 82.6 40.9 69.1 67.9
Contriever 77.2 67.1 76.7 75.2 65.1 53.7 75.4 79.2 83.3 39.4 69.2 68.3
GTE 77.4 62.6 74.7 75.8 62.0 51.8 77.8 80.0 75.0 39.4 67.6 67.2
[…]
Image
DSEwiki-ss 84.0 80.2 78.7 87.0 75.7 73.0 82.0 77.3 88.3 58.4 78.5 79.2
DSEdocmatix 82.1 77.2 79.6 87.8 73.9 72.4 81.7 83.1 89.4 67.9 79.5 80.1
ColPali 84.6 79.3 82.3 89.0 79.8 72.1 86.7 84.9 92.4 56.9 80.8 81.0
Table 5: Main results for page-level retrieval, with the best results in boldface and second best results underlined. For clarity, we omit results using VLM-text (Refer to Table A for full results).
5 Experiment
{tabu}
lll|cccccccccc|cc
Report &Indu. Worksh. Paper ure Report book ment Macro Micro
Recall@$k=1$
VLM-text
DPR 11.6 9.5 19.2 19.2 14.9 15.9 15.8 25.6 34.7 27.0 19.3 19.2
ColBERT 22.0 14.9 28.0 28.3 17.9 29.7 21.1 52.6 54.5 44.5 31.3 31.4
BGE 19.2 15.2 24.6 28.7 12.8 27.6 19.7 47.0 52.3 35.8 28.3 29.0
E5 15.9 8.8 27.7 24.3 14.6 21.8 14.7 45.6 53.0 40.5 26.7 26.4
Contriever 23.4 7.5 28.2 26.8 17.1 25.7 16.1 43.6 51.5 42.9 28.3 28.9
GTE 17.5 10.5 23.0 27.2 14.5 26.3 14.4 39.8 49.2 38.3 26.1 27.1
[…]
Image
DSEwiki-ss 20.6 15.1 31.0 31.1 20.1 29.2 22.0 39.3 37.5 35.8 28.2 29.2
DSEdocmatix 19.9 11.4 31.5 30.1 17.8 30.0 20.8 46.5 39.4 31.4 27.9 29.1
ColPali 22.5 21.3 36.6 30.9 26.8 32.1 19.3 52.5 51.8 33.6 32.7 32.5
Recall@$k=5$
VLM-text
DPR 31.0 25.7 36.7 44.9 33.0 34.1 34.9 49.9 56.3 51.1 39.8 40.4
ColBERT 41.8 37.7 53.7 61.8 35.1 52.4 46.1 83.2 70.1 62.5 54.4 56.0
BGE 41.0 28.1 52.7 59.2 36.7 46.0 50.7 72.0 71.5 59.9 51.8 53.2
E5 35.4 28.1 51.7 58.5 33.2 41.2 40.2 79.7 77.9 64.6 51.1 51.8
Contriever 40.2 29.2 54.1 57.9 36.8 47.1 44.6 68.8 76.2 61.9 51.7 53.0
GTE 36.7 25.6 51.2 56.6 39.7 46.6 46.4 72.2 74.3 63.2 51.3 52.3
[…]
Image
DSEwiki-ss 42.4 32.9 56.3 58.5 39.8 50.6 41.6 68.6 60.9 50.0 50.2 52.1
DSEdocmatix 39.6 36.2 53.9 57.5 33.7 52.5 42.8 69.4 63.1 48.9 49.8 51.9
ColPali 40.7 45.9 54.9 58.5 42.6 51.2 45.7 76.8 74.5 48.9 54.0 54.3
Recall@$k=10$
VLM-text
DPR 42.2 33.1 52.1 56.2 39.9 43.5 44.0 62.8 61.7 59.7 49.5 50.5
ColBERT 51.0 48.7 60.6 69.8 43.9 61.6 53.7 88.4 74.8 66.4 61.9 63.7
BGE 51.1 38.7 62.1 71.5 41.9 55.6 58.7 80.8 78.7 63.5 60.3 62.4
E5 45.3 38.6 62.0 70.5 45.6 50.0 55.3 87.1 82.4 66.8 60.4 61.2
Contriever 49.9 41.3 62.0 70.5 44.8 56.5 54.5 81.3 78.0 64.9 60.4 62.2
GTE 48.6 41.1 61.5 68.8 44.3 56.9 58.0 83.0 77.5 66.9 60.7 62.2
[…]
Image
DSEwiki-ss 55.9 41.3 61.5 68.1 47.8 60.7 54.2 72.9 68.3 54.4 58.5 61.1
DSEdocmatix 53.7 43.3 59.6 66.5 44.7 59.1 50.3 75.4 69.2 53.7 57.5 59.9
ColPali 53.6 54.1 64.4 69.5 48.8 60.7 54.0 81.9 82.5 50.4 62.0 63.2
Table 6: Main results for layout-level retrieval (Refer to Table A for full results with VLM-text and Hybrid inputs).
5.1 Evaluation Metric
The retriever scores each page or layout in the document based on its relevance to the question, and returns the top $k$ candidates with the highest scores. Recall@$k$ is defined as the proportion of ground truth page/layout evidence successfully retrieved. For layout matching, we calculate recall based on the overlaps between the bounding boxes of retrieved layouts and gold-standard layouts.
5.2 Baseline Models and Setting
We evaluate 6 state-of-the-art text retrievers: namely DPR, ColBERT, BGE, E5, Contriever, and GTE (see Appendix D.1). Additionally, we evaluate 5 VLM-based retrievers: 3 off-the-shelf models, namely DSEwiki-ss, DSEdocmatix, and ColPali (see Appendix D.2), and 2 models trained using MMDocIR training set (see Appendix C). Among all retrievers, ColBERT, ColPali, and Col-Phi3ours represent query/document as a list of token-level embeddings, while the other retrievers represent query/document as a single dense embedding. All retrievers are adapted to a dual-task setting:
  • Page Retrieval: For text retrievers, we use the text from OCR-page or VLM page as described in Section 3.4. For visual retrievers, we directly utilize document page screenshots.
  • Layout Retrieval: Text retrievers process multimodal layouts using OCR or VLM text (see Section 3.4). Visual retrievers process textual layouts using either Image input (cropped image of textual area) or Hybrid input (original text, as VLM can directly encode text).
5.3 Main Results for Page-level Retrieval
Table 4.2 presents the main results for page-level retrieval. Our key findings are as follows:
  • Superiority of Visual Retrievers: Visual retrievers consistently outperform text retrievers across various domains and retrieval metrics, highlighting the advantage of using screenshots to retain multimodal cues often lost in text conversion.
  • Effectiveness of MMDocIR: The visual retrievers trained on the MMDocIR training set demonstrate superior performance, demonstrating the value of high-quality training data.
  • Necessity of Token-level Embeddings: Compared to dense-level retrievers (e.g., BGE, DSE, DPR-Phi3ours), token-level retrievers (e.g., ColBERT, ColPali, Col-Phi3ours) achieve more advantageous results in Recall@1 and have marginal performance improvement in Recall@3/5. However, token-level embedding can incur storage costs of 10 times more than a single embedding (DSE requires 0.24GB for indexing MMDocIR while ColPali requires 10.0GB).
  • Top 5 Coverage: Retrieving top 5 pages provides substantial coverage of relevant information.
5.4 Main Results for Layout-level Retrieval
Table 5 shows the main results for layout-level retrieval. Our key findings are as follows:
  • Effectiveness of VLM-Text: Interestingly, VLM-text approaches perform comparably to visual retrievers, demonstrating the promising image description capabilities of state-of-the-art VLM. This greatly benefits textual retrievers in multimodal understanding.
  • Necessity of Token-level Embeddings: For layout retrieval tasks, token-level retrievers marginally outperform dense-level retrievers, demonstrating its importance of in such task.
  • Top 10 Coverage: For layout retrieval tasks, retrieving top 10 layouts does not guarantee comprehensive coverage of the ground truth layout labels, emphasizing the complexity of the tasks.
{tabu}
ll|ccc|ccc
Method   Page recall   Layout recall
OCR VLM $\Delta$ OCR VLM $\Delta$
$k=1$
DPR 22.3 27.2 +4.9 12.6 19.3 +6.7
ColBERT 40.3 45.8 +5.5 19.8 31.3 +11.5
BGE 35.7 40.6 +4.9 19.0 28.3 +9.3
E5 35.0 40.8 +5.8 18.4 26.7 +8.3
Contriever 35.3 40.9 +5.6 18.8 28.3 +9.5
GTE 35.4 38.9 +3.5 18.2 26.1 +7.9
$k=3\text{ or }5$
DPR 39.4 46.3 +6.9 23.7 39.8 +16.1
ColBERT 58.8 64.9 +6.1 33.2 54.4 +21.2
BGE 55.4 59.7 +4.3 32.7 51.8 +19.1
E5 54.8 60.3 +5.5 33.3 51.1 +17.8
Contriever 54.9 60.6 +5.7 31.7 51.7 +20.0
GTE 54.9 58.7 +3.8 33.5 51.3 +17.8
$k=5\text{ or }10$
DPR 49.0 57.8 +8.8 29.9 49.5 +19.6
ColBERT 66.0 72.3 +6.3 37.6 61.9 +24.3
BGE 62.7 68.4 +5.7 37.8 60.3 +22.5
E5 64.1 69.1 +5.0 39.0 60.4 +21.4
Contriever 63.1 69.2 +6.1 37.3 60.4 +23.1
GTE 63.2 67.6 +4.4 40.9 60.7 +19.8
Table 7: Results of text retrievers using OCR/VLM-text.
5.5 Text Retrieval: OCR-text vs VLM-Text
Text retrievers leveraging VLM-text significantly outperform those using OCR-text in both tasks. Based on results, OCR-text is insufficient for multimodal retrieval, while VLM-text retains richer multimodal information. Although VLM-text offers much more comprehensive text information than OCR-text, it also introduces higher computational overhead and longer inference time.
Most text retrievers based on on BERT Devlin et al. (2019), truncate input that exceed 512 tokens (approximately 380 english words). As shown in Figure 3(c), there are many pages containing more than 380 words (62.9% for OCR-page and 61.1% for VLM-page). Those pages suffer from critical information loss during page retrieval if the ground truth evidence is in the truncated part. In contrast, only a small fraction of layouts contain more than 380 tokens (3.9% for OCR-text, 4.8% for VLM-text, 0.5% for natural-text). Hence, as reflected in Table 4.2 and 5, text retriever demonstrates stronger performance on layout-level retrieval than on page-level retrieval.
{tabu}
ll|ccc
Method   Layout recall
Hybrid Image $\Delta$
$k=1$
DSEwiki-ss 24.6 28.2 +3.6
DSEdocmatix 27.5 27.9 +0.4
ColPali 28.5 32.7 +4.2
DPR-Phi3ours 28.9 29.5 +0.6
Col-Phi3ours 29.8 31.1 +1.3
$k=5$
DSEwiki-ss 46.7 50.2 +3.5
DSEdocmatix 48.2 49.8 +1.6
ColPali 52.2 54.0 +1.8
DPR-Phi3ours 50.1 51.6 +1.5
Col-Phi3ours 50.0 52.3 +2.3
$k=10$
DSEwiki-ss 55.8 58.5 +2.7
DSEdocmatix 57.4 57.5 +0.1
ColPali 60.0 62.0 +2.0
DPR-Phi3ours 55.5 60.2 +4.7
Col-Phi3ours 58.7 61.1 +2.4
Table 8: Results of visual retrievers: image vs hybrid.
5.6 Visual Retrieval: Image vs Hybrid input
Visual retrievers tend to perform better when encoding text as images via visual encoders, rather than processing native textual input with LLM backbones. This advantage largely stems from their training setup: visual retrievers are typically optimized using text queries paired with image-based passages or documents, but are not fine-tuned directly on purely textual passages. However, encoding text as images incurs substantial computational overhead. Representing text as image tokens requires significantly more resources than native text encoding. To address this inefficiency and promote balanced retrieval capabilities, we advocate for future visual retrievers to be jointly trained or fine-tuned on both text and visual retrieval tasks. Such hybrid training would enable models to efficiently process text when appropriate, without compromising performance on visual inputs.
{tabu}
l|ll|ccc Model Store Index Search
(GB) (MM:SS) (MM:SS)
Page Processing
Text
DPR 0.06 6:53 00:02
ColBERT 3.45 14:12 00:04
BGE 0.08 7:31 00:03
E5 0.08 8:26 00:03
[…]
[.2pt on 1pt off 1.5pt]2-6
Image
DPR-Phi3 0.24 101:20 00:04
ColPali 10.00 47:14 00:05
Col-Phi3 24.56 106:23 00:07
Layout Processing
Text
DPR 0.51 41:33 00:15
ColBERT 26.72 94:56 02:25
BGE 0.66 55:05 00:18
E5 0.66 60:21 00:18
[…]
[.2pt on 1pt off 1.5pt]2-6
Image
DPR-Phi3 1.99 735:51 00:44
ColPali 83.50 262:29 09:06
Col-Phi3 204.32 784:07 10:56
[…]
Hybrid
DPR-Phi3 1.84 130:38 01:09
ColPali 12.06 73:50 04:24
Col-Phi3 22.72 140:44 02:41
Table 9: Efficiency analysis different retrievers.
5.7 Efficiency Analysis
We evaluate the inference efficiency by measuring three key metrics : storage consumption, indexing time and, retrieval latency, as shown in Table  5.6. Experiments are conducted on a NVIDIA A100 40GB GPU with batch size of 4 for image and 256 for text. DPR-styled retrievers which generate single vector embeddings, demonstrates higher efficiency and lower computation across all metrics, compared to ColBERT-styled retrievers that produce token-level embeddings. Although DPR-styled retrievers slightly underperform in retrieval accuracy, their smaller embeddings size provide a significant advantage in the inference stage when storage space and inference time are concerned.
Another key finding is that textual inputs are significantly more efficient than the visual inputs across all metrics. Meanwhile, hybrid retrieval system, which processes text in the image through LLM rather than visual encoders, further reduces memory and time consumption. Hence, future works on training hybrid retrieval system are encouraged as it offers a strong balance between computational efficiency and retrieval performance.
6 Related Work
DocCVQA Tito et al. (2021) proposes extracting information from a document image collection. However, it is limited by its small question set (20 questions). While PDF-MVQA Ding et al. (2024) is tailored for multimodal retrieval in biomedical articles, it is annotated by GPT-3.5-turbo rather than experts. SciMMIR Wu et al. (2024) also investigates multimodal retrieval but only provides image-caption pairs, lacking user queries paired with the corresponding document pages. Ma et al. (2024a) introduces two relevant datasets, namely Wiki-SS and DocMatix-IR. Wiki-SS is derived from natural questions Kwiatkowski et al. (2019) , wherein evidence passages are screenshots of Wikipedia pages. However, natural questions are primarily designed for text retrieval, and the provided screenshots may not consistently capture the ground-truth evidence, as only the front page is considered. DocMatix-IR is constructed from the large-scale DocMatix Laurençon et al. (2024) dataset using filtering and hard negative mining. However, the questions are generated by Phi-3-small Abdin et al. (2024) rather than human experts, and are not de-contextualized for retrieval task. ViDoRe Faysse et al. (2024) is the most relevant benchmark to MMDocIR. It integrates multiple DocVQA datasets and provides new industrial documents. Upon a thorough examination of the 2,400 questions, we find that over 80% questions exhibit notable limitations in terms of their complexity, contextual clarity, and the absence of complete document corpora. Refer to Appendix F for detailed quantitative and qualitative analysis of ViDoRe.
7 Conclusion
In conclusion, multimodal document retrieval presents a complex challenge that requires both understanding and integrating diverse data modalities beyond plain text. To more effectively evaluate these capabilities, we introduce the MMDocIR benchmark, which features the innovative dual-task retrieval capabilities targeting page-level and layout-level document granularity. The MMDocIR includes a rich dataset featuring expertly annotated labels for 1,685 questions and bootstrapped labels for 73,843 questions, serving as a valuable resource for both training and evaluation of multimodal document retrieval. Our comprehensive empirical studies show that visual-driven retrievers significantly outperform text-driven ones, underscoring the importance of visual information in improving retrieval performance. Future work can expand upon these findings by optimizing retrieval algorithms to enhance both accuracy and efficiency of multimodal document retrieval systems.
Limitations
The limitations of MMDocIR are summarized as follows:
  • Incomplete layout label annotations for training set: For 3 out of 7 training subsets, our semi-automated pipelines could not extract layout labels. These pipelines are optimized for datasets with single text or image layouts and cannot handle complex or cross-modal layouts. Future work should explore leveraging advanced vision-language models (VLMs) to facilitate annotation of layout labels for these subsets.
  • Lack of joint text and visual training: As demonstrated in Section 5.6, all visual retrievers are suboptimal at modeling text passages, compared to modeling text as image screenshots. Our current visual retrievers do not explicitly utilize text query-document pairs to address this limitation. Future research should consider integrating both text and visual passages for joint training or finetuning to improve performance on both retrieval tasks.
Appendix Table of Contents
Due to page limits, we provide meaningful descriptions, results, implementation details, case studies, and other supplementary materials in Appendix. Our 21-page appendix is organized as follows:
  • Appendix A: Supplementary experimental results for page and layout retrieval.
  • Appendix B: process and implementation for MMDocIR curation.
    • Related DocVQA benchmarks (§B.1)
    • Multimodal corpora selection (§B.2)
    • Question filtering for IR context (§B.3)
    • Collection of training multimodal documents (§B.4)
    • Curation of ground-truth training labels (§B.5)
    • Selection of hard negatives (§B.6)
    • Modality distribution (§B.7)
    • Resource and Artifact URLs (§B.8)
  • Appendix C: Training visual retrievers with MMDocIR training set.
    • Encoding text queries and images (§C.1)
    • Similarity calculation (§C.2)
    • Training objectives (§C.3)
    • Implementation details (§C.4)
  • Appendix D: Text and visual retriever details.
    • Text-centric retrieval (§D.1)
    • Vision-driven retrieval (§D.2)
    • Retriever implementation (§D.3)
  • Appendix E: Examples from MMDocIR.
    • Document pages (10 domains) (§E.1)
    • Document layouts (§E.2)
    • Page and layout annotations (§E.3)
  • Appendix F: Analyses on ViDoRe benchmark.
    • Search query and new annotation (§F.1)
    • Corpora analysis (§F.2)
  • Appendix G: License agreements.
  • Appendix H: Ethical considerations.
Appendix A Supplementary Experimental Results
In this section, we provide the full results of page-level and layout-level retrievals, which supplement our main results discussion in Section 5.3 and Section 5.4 respectively.
Specifically, Table A extends the main page-level results shown in Table 4.2 with the results of text retrievers using OCR-text. Table A extends the main layout-level results shown in Table 5 with the results of (i) text retrievers using OCR-text and (ii) visual retrievers using hybrid inputs.
{tabu}
lll|cccccccccc|cc